OpenStudy (anonymous):

how to solve cot[arcsin(-sqrt2/9)]

OpenStudy (anonymous):

cot=cos/sin sin^2+cos^2=1 cos=sqrt(1-sin^2) cot=sqrt(1-sin^2)/sin

OpenStudy (tkhunny):

1) Please draw a right triangle. 2) Designate an angle. 3) Label two sides so that the sine of that angle is \(\sqrt{2}/9\) 4) Use the Pythagorean Theorem to find the length of the third side. 5) The cotangent should now be trivial - almost. 6) The original sine was negative. This is Quadrants III and IV. You will have to figure out the correct sign of the cotangent you seek.